# ReMoDetect: Reward Models Recognize Aligned LLM’s Generations

### v1

@@ Abstract
The remarkable capabilities and easy accessibility of large language models (LLMs) have significantly increased societal risks (e.g., fake news generation), necessitating the development of LLM-generated text (LGT) detection methods for safe usage. However, detecting LGTs is challenging due to the vast number of LLMs, making it impractical to account for each LLM individually; hence, it is crucial to identify the common characteristics shared by these models. In this paper, we draw attention to a common feature of recent powerful LLMs, namely the alignment training, i.e., training LLMs to generate human-preferable texts. Our key finding is that as these aligned LLMs are trained to maximize the human preferences, they generate texts with higher estimated preferences even than human-written texts; thus, such texts are easily detected by using the reward model (i.e., an LLM trained to model human preference distribution). Based on this finding, we propose two training schemes to further improve the detection ability of the reward model, namely (i) continual preference fine-tuning to make the reward model prefer aligned LGTs even further and (ii) reward modeling of Human/LLM mixed texts (a rephrased texts from human-written texts using aligned LLMs), which serves as a median preference text corpus between LGTs and human-written texts to learn the decision boundary better. We provide an extensive evaluation by considering six text domains across twelve aligned LLMs, where our method demonstrates state-of-the-art results. Code is available at https://github.com/hyunseoklee-ai/reward_llm_detect.
@@ 3.3 Reward Modeling of Human and LLM Mixed Dataset
Detection stage. After training ReMoDetect, we use the predicted reward score $r_{\phi}(x,y)$ to determine whether the given text is LGT or human-written texts where a higher score indicates LGT. Unlike recent detection schemes that require multiple forwards (for perturbing the input [10, 12]), ReMoDetect only requires a single forward pass, thus showing inference efficiency (in Appendix B.4).
@@ 4 Experiments
Evaluation setup. We mainly report the area under the receiver operating characteristic curve (AUROC) as a threshold-free evaluation metric (results with other metrics are presented in Appendix B.7). Here, the text is written (or generated) in 6 text domains introduced in Fast-DetectGPT [12] and MGTBench [14], including PubMed [28], XSum [34], Reuters [42], Essay [42], and WritingPrompts [4] (each benchmark consists of different types of WritingPrompts, thus denoting the version in [12] as small-sized). In addition to GPT3.5 Turbo, GPT4, and Claude, which are already provided in the benchmark, we consider more aligned LLMs $\mathcal{M}$, including Llama3 70B instruct [40], Claude3 Opus [5] Gemini pro [37], and GPT4 Turbo [25]. We also consider more aligned LLMs, e.g., models trained with direct preference optimization (DPO) [29], in Appendices B.2 & B.3.
@@ B.2 Comparison on Additional Aligned LLMs
B.3 Comparison on Aligned LLMs trained with Direct Preference Optimization
Table 9: AUROC(%) on Fast-DetectGPT benchmark [12] under aligned LLM trained with DPO. Phi-3 mini, small, and medium have a parameter size of 3.8B, 7B, and 14B, respectively.
Model Domain Loglik. Rank D-GPT LRR NPR FD-GPT Open-D Chat-D Ours
Phi-3 mini PubMed 65.0 56.2 46.8 48.8 45.4 63.7 37.7 80.7 94.5
XSum 70.3 64.1 69.0 61.7 70.5 91.0 82.7 23.4 97.6
WP-s 82.4 73.3 89.6 68.8 87.8 96.7 60.0 31.1 99.3
Phi-3 small PubMed 57.2 50.4 48.1 51.2 48.1 59.9 31.9 82.7 91.7
XSum 81.1 69.7 70.0 68.8 72.7 95.6 79.3 19.5 98.7
WP-s 84.0 72.3 86.7 67.1 83.2 97.2 58.6 32.2 97.4
Phi-3 medium PubMed 65.4 55.4 51.2 50.3 37.6 61.7 34.2 15.8 95.2
XSum 64.5 61.2 80.7 79.0 81.3 85.4 75.0 18.1 98.0
WP-s 83.1 73.6 90.3 70.6 90.2 95.7 53.9 38.5 98.8
We additionally consider aligned LLMs that do not use reward models for alignment training. To this end, we consider aligned LLMs that use Direct Preference Optimization (DPO) [29], an alternative alignment training to Reinforcement Learning with Human Feedback (RLHF). Note that a recently released Phi-3 [24] only uses DPO (followed by supervised fine-tuning) for alignment training and shows remarkable performance in various domains, thus being considered an aligned LLM in our experiment. As shown in Table 9, ReMoDetect also outperforms baselines in all cases, showing that our method can be applicable even if aligned LLMs are not trained with reward models.
B.4 Inference time efficiency
Method Detection Time (secs) Model Parameters AUROC
Log-likelihood 0011.7 2.7B 88.6
OpenAI-D 0004.4 0.2B 68.9
DetectGPT 7738.8 3B & 2.7B 79.2
NPR 7837.3 3B & 2.7B 78.7
Fast-DetectGPT 0062.7 6B & 2.7B 91.9
Ours 0008.7 0.5B 95.8
In Table 10, we compared detection time, model parameter size, and average AUROC on the Fast-DetectGPT benchmark. The detection time was measured in an A6000 GPU, and the overall detection time was measured with 300 samples of the human/GPT3.5 Turbo XSum dataset. ReMoDetect shows the best average AUROC performance among the methods, but 7.2 times faster, and uses a 17.4 times smaller model than the second best model, Fast-DetectGPT. We see that ReMoDetect is a very effective detection method.
B.5 Dataset Examples and Predicted Reward Scores
Model Domain Example ReMo-D Score
In Table 11, we show dataset examples and their ReMoDetect score.
B.6 Additional Observational Studies
(a) Reuters
(b) Essay
(c) WritingPrompts
(d) Reuters
(e) Essay
(f) WritingPrompts
(a) Reuters
(b) Essay
(c) WritingPrompts
(d) Reuters
(e) Essay
(f) WritingPrompts
B.7 Additional Performance Metric
@@
B.8 Comparison with GPTZero per domain.
@@
Appendix C Societal Impact
This paper presents ReMoDetect that improves the performance of detecting aligned LGTs. We expect that our approach will show numerous positive impacts by detecting LGTs, such as in fake news and academic plagiarism. One possible negative impact can be the improved adversarial mechanism followed by the improved detection method (i.e., ReMoDetect); thus, incorporating such a scenario will be an interesting future direction to explore, where we believe using ReMoDetect to such a scenario can be promising (as it shows robustness in multiple cases in Section 4.3).

### v2

@@ Abstract
The remarkable capabilities and easy accessibility of large language models (LLMs) have significantly increased societal risks (e.g., fake news generation), necessitating the development of LLM-generated text (LGT) detection methods for safe usage. However, detecting LGTs is challenging due to the vast number of LLMs, making it impractical to account for each LLM individually; hence, it is crucial to identify the common characteristics shared by these models. In this paper, we draw attention to a common feature of recent powerful LLMs, namely the alignment training, i.e., training LLMs to generate human-preferable texts. Our key finding is that as these aligned LLMs are trained to maximize the human preferences, they generate texts with higher estimated preferences even than human-written texts; thus, such texts are easily detected by using the reward model (i.e., an LLM trained to model human preference distribution). Based on this finding, we propose two training schemes to further improve the detection ability of the reward model, namely (i) continual preference fine-tuning to make the reward model prefer aligned LGTs even further and (ii) reward modeling of Human/LLM mixed texts (a rephrased texts from human-written texts using aligned LLMs), which serves as a median preference text corpus between LGTs and human-written texts to learn the decision boundary better. We provide an extensive evaluation by considering six text domains across twelve aligned LLMs, where our method demonstrates state-of-the-art results. Code is available at https://github.com/hyunseoklee-ai/ReMoDetect.
@@ 3.3 Reward Modeling of Human and LLM Mixed Dataset
Detection stage. After training ReMoDetect, we use the predicted reward score $r_{\phi}(x,y)$ to determine whether the given text is LGT or human-written texts where a higher score indicates LGT. Unlike recent detection schemes that require multiple forwards (for perturbing the input [11, 13]), ReMoDetect only requires a single forward pass, thus showing inference efficiency (in Section 4.3).
@@ 4 Experiments
Evaluation setup. We mainly report the area under the receiver operating characteristic curve (AUROC) as a threshold-free evaluation metric (results with other metrics are presented in Appendix B.3). Here, the text is written (or generated) in 6 text domains introduced in Fast-DetectGPT [13] and MGTBench [15], including PubMed [29], XSum [35], Reuters [43], Essay [43], and WritingPrompts [4] (each benchmark consists of different types of WritingPrompts, thus denoting the version in [13] as small-sized). In addition to GPT3.5 Turbo, GPT4, and Claude, which are already provided in the benchmark, we consider more aligned LLMs $\mathcal{M}$, including Llama3 70B instruct [41], Claude3 Opus [5] Gemini pro [38], and GPT4 Turbo [26]. We also consider more aligned LLMs, e.g., models trained with direct preference optimization (DPO) [30], in Table 6 and Appendix B.2.
@@ 4.3 Additional Analysis
Table 6: LGT Detection results on non-RLHF trained LLMs. We report AUROC (%) of multiple LGT detection methods, including log-likelihood (Loglik.), Rank, Fast-DetectGPT (FD-GPT), OpenAI-Detector (Open-D), ChatGPT-Detector (Chat-D), and ReMoDetect (Ours). We consider LGT detection benchmarks from Fast-DetectGPT: PubMed, XSum, and WritingPrompts-small (WP-s). Here, Phi-3 medium is DPO trained and OLMo-7B-SFT is SFT-only trained. The bold indicates the best result within the group.
Model Domain Loglik. Rank FD-GPT Open-D Chat-D Ours
Phi-3 mini PubMed 65.0 56.2 63.7 37.7 80.7 94.5
XSum 70.3 64.1 91.0 82.7 23.4 97.6
WP-s 82.4 73. 96.7 60.0 31.1 99.3
Phi-3 small PubMed 57.2 50.4 59.9 31.9 82.7 91.7
XSum 81.1 69.7 95.6 79.3 19.5 98.7
WP-s 84.0 72.3 97.2 58.6 32.2 97.4
Phi-3 medium PubMed 65.4 55.4 61.7 34.2 15.8 95.2
XSum 64.5 61.2 85.4 75.0 18.1 98.0
WP-s 83.1 73.6 95.7 53.9 38.5 98.8
OLMo 7B-SFT PubMed 88.4 60.5 92.8 62.0 23.6 94.1
XSum 96.6 66.0 99.1 97.3 05.9 98.1
WP-s 98.1 78.5 98.8 95.2 19.5 99.2
Average - 86.0 63.8 91.2 72.2 43.8 95.3
ReMoDetect for non-RLHF aligned LLMs. We additionally consider aligned LLMs that do not use reward models for alignment training, i.e., non-RLHF trained LLMs. To this end, we consider aligned LLMs that use Direct Preference Optimization (DPO) [30], an alternative alignment training to RLHF. Note that a recently released Phi-3 [25] only uses DPO (followed by supervised fine-tuning; SFT) for alignment training and shows remarkable performance in various domains, thus being considered an aligned LLM in our experiment. As shown in Table 6, ReMoDetect also outperforms baselines in all cases, showing that our method can be applicable even if aligned LLMs are not trained with reward models. Furthermore, we also considered the detection scenario for the SFT-only model that does not use the alignment training. Here, we observe that ReMoDetect effectively detects the LGTs from the SFT-only model as well as outperforming other baselines. We believe this is because the SFT implicitly trains the model to reflect the human preference from the instruction tuning dataset [10], thus making the ReMoDetect well-detect the texts from SFT models.
@@
Method Detection Time (secs) Model Parameters AUROC
Log-likelihood 0011.7 2.7B 88.6
DetectGPT 7738.8 3B & 2.7B 79.2
NPR 7837.3 3B & 2.7B 78.7
Fast-DetectGPT 0062.7 6B & 2.7B 91.9
Ours 0008.7 0.5B 95.8
Inference time efficiency. In Table 8, we compared detection time, model parameter size, and average AUROC on the Fast-DetectGPT benchmark. The detection time was measured in an A6000 GPU, and the overall detection time was measured with 300 samples of the human/GPT3.5 Turbo XSum dataset. ReMoDetect shows the best average AUROC performance among the methods, but 7.2 times faster, and uses a 17.4 times smaller model than the second best model, Fast-DetectGPT.
@@ 5 Discussion and Conclusion
Societal impact. This paper presents ReMoDetect that improves the performance of detecting aligned LGTs. We expect that our approach will show numerous positive impacts by detecting LGTs, such as in fake news and academic plagiarism. One possible negative impact can be the improved adversarial mechanism followed by the improved detection method (i.e., ReMoDetect); thus, incorporating such a scenario will be an interesting future direction to explore, where we believe using ReMoDetect to such a scenario can be promising (as it shows robustness in multiple cases in Section 4.3).
Acknowledgements
We thank Jongheon Jeong and Myungkyu Koo for providing helpful feedback and suggestions in preparing an earlier version of the manuscript. This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST)) and NIPA(National IT Industry Promotion Agency), through the Ministry of Science and ICT (Hyperscale AI flagship project).
@@ B.2 Comparison on Additional Aligned LLMs
B.3 Additional Performance Metric
@@
B.4 Dataset Examples and Predicted Reward Scores
Model Domain Example ReMo-D Score
In Table 12, we show dataset examples and their ReMoDetect score.
B.5 Additional Observational Studies
(a) Reuters
(b) Essay
(c) WritingPrompts
(d) Reuters
(e) Essay
(f) WritingPrompts
(a) Reuters
(b) Essay
(c) WritingPrompts
(d) Reuters
(e) Essay
(f) WritingPrompts
B.6 Robustness of Reward Models against Rephrasing Attacks
Table 13: Robustness against rephrasing attacks. We report the average AUROC (%) before (‘Original’) and after (‘Attacked’) the rephrasing attack with T5-3B on the Fast-DetectGPT benchmark, including XSum, PubMed, and WritingPrompts-small. Values in the parenthesis indicate the relative performance drop after the rephrasing attack. The bold indicates the best result.
Model Accuracy Loglik. D-GPT NPR FD-GPT Ours (reward model) Ours (ReMoDetect)
GPT4 Original 82.1 69.0 68.1 90.6 79.0 97.9
Attacked 63.7 (-22.4%) 44.8 (-35.1%) 47.0 (-31%) 74.5 (-17.7%) 71.2 (-9.9%) 87.2 (-10.9%)
Llama3 70B Original 93.5 84.9 84.9 96.8 80.9 98.5
Attacked 79.9 (-14.5%) 61.7 (-27.4%) 64.7 (-23.7%) 87.9 (-9.2%) 71 (-12.3%) 88.3 (-10.4%)
Gemini pro Original 79.2 71.3 75.8 79.9 64.1 81.8
Attacked 64.9 (-18%) 50.7 (-28.9%) 55.7 (-26.6%) 64.5 (-19.3%) 55.8 (-13%) 67.4 (-17.6%)
In Table 13, we compare the robustness against the paraphrased attack of the reward model and other baselines including ReMoDetect. The experiment shows that the reward model is robust against paraphrasing attacks (i.e. reward model and ReMoDetect are the two least drops against paraphrasing attacks). From the results, we hypothesize that the robustness against attack came from the reward model itself. Conceptually the human preference for the text samples doesn’t change much as the distribution shifts or paraphrases some words, hence, the reward score is independent of the minor variation of the sentence. We believe that the result of the experiment supports our hypothesis. Furthermore, exploring the characteristics and applications of the reward model would be interesting in the future.
B.7 Additional ReMoDetect Models Trained From Differently Initialized Reward Models.
Table 14: Comparison of multiple ReMoDetect models trained from reward models, including deberta, Gemma-2B (G. 2B), Llama3-8B (L. 8B). We report the average AUROC (%) on the fastdetectGPT benchmark, including PubMed, XSum, and WritingPrompts-small (WP-s).
Model Domain FD-GPT Open-D Ours (deberta) Ours (G. 2B) Ours (L. 8B)
GPT3.5 Turbo PubMed 090.2 61.9 096.4 090.1 094.7
XSum 099.1 91.5 099.8 100.0 100.0
WP-s 099.2 70.9 099.9 099.9 099.7
GPT4 PubMed 085.0 53.1 096.1 091.4 092.1
XSum 090.7 67.8 098.8 099.9 100.0
WP-s 096.1 50.7 098.7 099.6 099.4
GPT4 Turbo PubMed 88.8 55.8 097.0 091.2 092.9
XSum 97.4 88.2 099.8 100.0 100.0
WP-s 99.4 72.3 100.0 100.0 100.0
Llama3 70B PubMed 90.8 52.9 096.3 091.8 094.3
XSum 99.7 96.2 099.5 100.0 099.9
WP-s 99.9 77.5 099.8 099.6 099.6
Gemini pro PubMed 82.1 57.3 085.6 078.8 081.8
XSum 79.5 72.2 088.2 087.5 085.3
WP-s 78.0 70.2 071.6 084.2 089.2
Calude3 Opus PubMed 88.2 48.9 096.4 090.9 093.3
XSum 96.2 86.2 099.5 099.9 099.8
WP-s 93.5 65.7 099.9 099.7 099.8
Average - 91.9 68.9 095.8 094.6 095.6
We additionally consider the ReMoDetect models trained from differently initialized reward models. To address the consideration, we conducted experiments to train ReMoDetect using three reward models. As shown in Table 14, ReMoDetect models consistently outperform other baselines, even though the model trained from differently initialized reward models. Nonetheless, the ReMoDetect’s detection performance can vary with initialization. Thus, we suggest interesting future works to find a better detector, such as ensembling several trained models or using an enhanced reward model.
B.8 Comparison with GPTZero per domain
@@
B.9 Comparison on Aligned Small Language Models
Table 16: AUROC (%) of multiple LGT detection methods, including log-likelihood (Loglik.), Rank, Fast-DetectGPT (FD-GPT), OpenAI-Detector (Open-D), ChatGPT-Detector (Chat-D), and ReMoDetect (Ours). We consider LGT detection benchmarks from Fast-DetectGPT: PubMed, XSum, and WritingPrompts-small(WP-s). The bold indicates the best result within the group.
Model Domain Loglik. Rank FD-GPT Open-D Chat-D Ours
Llama3 8B-it PubMed 85.0 60.4 89.6 53.7 33.4 94.6
XSum 82.3 68.9 86.8 95.4 13.1 85.4
WP-s 87.2 72.3 91.0 81.2 26.4 95.5
Gemma2 9B-it PubMed 69.8 55.9 71.6 36.4 85.1 95.1
XSum 85.1 69.4 94.0 74.0 97.7 99.5
WP-s 86.7 71.9 96.6 50.1 70.3 96.8
Gemma2 2B-it PubMed 67.9 56.6 72.3 44.4 78.1 90.0
XSum 82.1 18.2 89.8 67.6 97.2 94.9
WP-s 84.6 71.8 99.0 70.8 63.7 94.2
Qwen2 1.5B-it PubMed 82.3 61.0 89.8 62.9 23.9 92.7
XSum 96.5 66.7 98.3 97.2 01.3 99.6
WP-s 97.5 78.2 98.6 94.3 17.7 99.1
OLMo 7B-sft PubMed 88.4 60.5 92.8 62.0 23.6 94.1
XSum 96.6 66.0 99.1 97.3 05.9 98.1
WP-s 98.1 78.5 98.8 95.2 19.5 99.2
Average - 86.0 63.8 91.2 72.2 43.8 95.3
We additionally consider small aligned models particularly when the model parameter size is smaller than 10B, including Llama3-8b, Gemma-2-9b, Gemma-2-2b, Qwen2-1.5b-it, and Olmo7b-sft. As shown in Table 16, ReMoDetect also effectively detects LGT of small language models. For instance, ReMoDetect achieves 97.1% average AUROC in Qwen2-1.5b-it while the second-best reaches 84.8%.